OpenStudy (anonymous):

Help with Precal 3x^4+2x^2-6x+1 all over x-1

OpenStudy (raden):

use the long division or use the horner's diagram

OpenStudy (anonymous):

Can you give an example? i get confused with long division.

OpenStudy (raden):

it like you dividing 17 by 3 :v

ganeshie8 (ganeshie8):

familiar with synthetic division ?

ganeshie8 (ganeshie8):

\( \begin{array}{} 1~|&3&0&2&-6&1\\ &&&&\\ \hline &&&&\\ &3&&& \end{array} \)

OpenStudy (anonymous):

Oh yeah! So it will be 1 x 3 = 3 . then do i do 3x0 or 3x1?

OpenStudy (anonymous):

sorry i was away from the computer for a while, Im here now :)

ganeshie8 (ganeshie8):

\( \begin{array}{} 1~|&3&0&2&-6&1\\ &&3&&\\ \hline &&&&\\ &3&&& \end{array} \)

ganeshie8 (ganeshie8):

next, add 3+0 and put it down in the last row

ganeshie8 (ganeshie8):

\( \begin{array}{} 1~|&3&0&2&-6&1\\ &&3&&\\ \hline &&&&\\ &3&3&& \end{array} \)

OpenStudy (anonymous):

Oh crap right! so it will then be 3 3 on bottom row?

OpenStudy (anonymous):

lol gotcha. then i do 3 +2?

ganeshie8 (ganeshie8):

yes :) next, multiply 3*1 and put it in next column in middle row : \( \begin{array}{} 1~|&3&0&2&-6&1\\ &&3&3&\\ \hline &&&&\\ &3&3&& \end{array} \)

OpenStudy (anonymous):

I tried to type mine like yours using the equation button but I cant figure it out :P

ganeshie8 (ganeshie8):

you got it :)

OpenStudy (anonymous):

so 3 3 5

OpenStudy (anonymous):

3 3 5 -1 0?

ganeshie8 (ganeshie8):

\( \begin{array}{} 1~|&3&0&2&-6&1\\ &&3&3&~~5&-1\\ \hline &&&&\\ &3&3&5&-1&|~0 \end{array} \)

ganeshie8 (ganeshie8):

Correct ! so 0 is remainder and do u knw how to figure out quotient ha ?

OpenStudy (anonymous):

hang on let me find my notes! lol

ganeshie8 (ganeshie8):

okay ! u can figure out quotient from last row

OpenStudy (anonymous):

its 0?

ganeshie8 (ganeshie8):

last row has below : \(3~~3~~5~~-1~~|0 \) so, the remainder is : \(0\) quotient is : \(3x^2 + 3x^2 +5x - 1\)

ganeshie8 (ganeshie8):

quotient is ur answer btw :)

ganeshie8 (ganeshie8):

\(\large \frac{3x^4+2x^2-6x+1}{x-1} = 3x^3 + 3x^2 + 5x-1\)

OpenStudy (anonymous):

Oh my lol. I knew it was simple just getting there I dont want to make a mistake you know?

OpenStudy (anonymous):

Can I try the next one and you check it, if you have time?

ganeshie8 (ganeshie8):

sure :)

ganeshie8 (ganeshie8):

corrected the typo : quotient is : \(3x^{\color{Red}{3}} + 3x^2 +5x - 1 \)

OpenStudy (anonymous):

so 3x^3?

OpenStudy (anonymous):

5x^4+5x^2+5 all over x^2-x+1

ganeshie8 (ganeshie8):

yes, 3x^3 + 3x^2 + 5x - 1

OpenStudy (anonymous):

how did you type it all fancy?

ganeshie8 (ganeshie8):

its latex :)

ganeshie8 (ganeshie8):

copy paste below : ``` \( \color{red}{heyyyyyyyyyy} \) ```

OpenStudy (anonymous):